Consider the set of equations


5x1 − 4x2 + 3x3 + x4 = 3

2x1 + x2 + 5x3 − 3x4 = 0

x1 + 6x2 − 4x3 + 2x4 = 15


where x1 = 1, x2 = 2, x3 = 1, x4 = 3 is a feasible solution. Is this solution feasible solution

If not, reduce this feasible solution to two different basic feasible solution.


1
Expert's answer
2021-09-10T10:10:30-0400

"5x_1 − 4x_2 + 3x_3 + x_4 = 3...(1)

\\2x_1 + x_2 + 5x_3 − 3x_4 = 0 ...(2)

\\x_1 + 6x_2 − 4x_3 + 2x_4 = 15 ...(3)"

Put "x_1 = 1, x_2 = 2, x_3 = 1, x_4 = 3" in (1), we get

"5\times 1-4\times2+3\times1+3=3
\\\Rightarrow 3=3"

Put "x_1 = 1, x_2 = 2, x_3 = 1, x_4 = 3" in (2), we get

"2\times1+2+5\times1-3\times3=0
\\\Rightarrow 0=0"

Put "x_1 = 1, x_2 = 2, x_3 = 1, x_4 = 3" in (3), we get

"1+6\times2-4\times1+2\times3=15
\\\Rightarrow 15=15"

"\therefore" "x_1 = 1, x_2 = 2, x_3 = 1, x_4 = 3" is a feasible solution.